OpenStudy (anonymous):

given that f(x)=x^2+3x-10 and g(x)=4x+3 find (f o g)(x)

OpenStudy (mathmate):

f(4x+3)=(4x+3)^2+3(4x+3)-10 =16x^2+24x+9 +12x+9 -10 =16x^2 + 36x +8

OpenStudy (anonymous):

huh?

OpenStudy (mathmate):

fog(x)=f(g(x))

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

what if it was gof(1)?

OpenStudy (anonymous):

then it is g(f(1))